Fermi National Laboratory
 

Operation Summary for February 10, 2010 - February 12, 2010

 

Wednesday Feb. 10

The day shift began with interlock technicians conducting Tevatron (TeV) safety system tests, with Operations monitoring the Recycler stash (242.3E10), with the Antiproton Source stacking (24.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store.

During the day shift, Operations reported that MTest took beam until the setup for a new experiment (T944 Jasmin) prompted an access at 8:34 AM. Accesses and safety system test continued in the TeV and collision halls until 3 PM. TeV turn-on and checkout followed.

During the evening shift, Operations disabled the MTest beam at 6 PM. Operations put the TeV into HEP shot setup at 9:12 PM and established store 7595 at 10:21 PM with an Average Initial Luminosity (AIL) of 270.2E30.

Thursday Feb. 11

The midnight shift began with Operations monitoring store 7595 (176.77E30) and the stash (179.8E10), with Pbar stacking (15.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to MTest at 4:12 AM.

During the day shift, Operations terminated store 7595 at 9:45 AM and put the TeV into shot setup. Operations established store 7597 at 11:10 AM with an AIL of 186.1E30. At 3:27 PM, store 7597 quenched at sector A3; quench recovery and investigation followed.

During the evening shift, Operations disabled the MTest beam at 6 PM. At 6:58 PM, a Pbar Accumulator quadrupole power supply (A:LQ) tripped off. Operations called in EE Support personnel and they had the supply back on by 10:48 PM. Operations established store 7601 at 11:11 PM with an AIL of 188.75E30.

Friday Feb. 12

The midnight shift began with Operations monitoring store 7601 (141.15E30) and the stash (22.0E10), with Pbar stacking (21.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to MTest at 4 AM. Store 7601 aborted at 4:40 AM due to a bad CAMAC crate power supply in the TeV. Operations put the TeV through a dry squeeze, a shot setup, and established store 7603 at 7:05 AM with an AIL of 154.3E30.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/12/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory